OpenStudy (anonymous):

A wire 5.1 m long experiences a magnetic force of 0.61 N due to a perpendicular uniform magnetic field. If the wire carries a current of 10.0 A, what is the magnitude of the magnetic field? answer in units of T.

OpenStudy (espex):

\[F =\beta I l\]\[\beta =\frac{F}{Il}\]\[\frac{0.61N}{10A \times 5.1m}=0.012\frac{N}{Am}\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (espex):

You are welcome.